# General Place Recognition Survey: Towards Real-World Autonomy

### v1

@@ IV-D Efficiency
Although sequence matching improves matching accuracy, the brute-force method used in SeqSLAM [78] is time-consuming. Including odometry and movement information may help enhance performance. Various approaches have been proposed to boost SeqSLAM’s efficiency, such as particle filters [96], approximate nearest neighbor searches [97], and Hidden Markov Models [98]. However, these methods largely rely on environmental conditions, limiting their robustness in dynamic and challenging settings. Recent works [121] and [111] have introduced frameworks that balance efficiency and accuracy by refining SeqSLAM with a coarse-to-fine search strategy. Similarly, in several SLAM systems, odometry information can constrain the search space for PR, thereby providing a starting point for the place node identification.
@@ V-A Long-Term & Large-Scale Navigation
The most in-demand robotic tasks requiring PR is the autonomous navigation, including applications such as autonomous driving [128] and subterranean search [129]. PR enables robots to obtain their global location with precision up to the topological level in a known environment, despite condition changes of environments. This capability is crucial to safe and reliable navigation, as it allows robots to (1) determine whether this place is visited before, (2) recorver from failure against the kidnapped problem, and (3) progressively update and enhance their navigational maps over time.
@@ V-A1 Brief Survey
Large-scale SLAM often benefits from PR solutions that enable robots to recognize previously visited locations against significant environmental changes or different viewpoints, thereby correcting drift errors and enhancing map accuracy. Typical SLAM systems such as ORB-SLAM [130] and VINS-Mono [131] are designed with the PR module to maintain the global consistency of poses, mainly rely on the DBoW2 library [115] and use the BoW representation with specific global descriptors to quickly detect loop candidates. The additional RANSAC-based verification algorithms is also coupled to reject wrong place candidates. LiDAR- and Radar-based SLAM [132, 133] also benefits from the development of LPR and RPR approaches, where the need for an extra camera to find a loop is eliminated.
@@
Tracking failure, which often refers to a situation where the system fails to correctly associate across adjacent frames, is the primary cause to make SLAM collapse and induce the kidnapped problem. It may encouter in practice due to reasons such as motion blur, occlusion, and hardware disconnection. Recovering from failure by relocalizing the robot in the previously built map necessitates solving a more complex PR problem than loop detection since odometry prior is unavailable. Chen et al. [134] developed a submap-based SLAM system that enhances resilience and mapping integrity by initiating a new process to manage a submap upon failure, which is integrated into the main map after detecting a loop closure using DBoW2 to ensure continuous and accurate mapping. Kuse and Shen [135] streamlined the training of NetVLAD [45] for VPR by introducing an all-pair loss function and decoupled convolutions. This innovation leads to quicker training convergence and reduced the number of learnable parameters. And the enhanced NetVLAD method was effectively integrated into a stereo-inertial SLAM system, achieving real-time loop closure detection and failure recovery in complex indoor settings.
@@
Global localization broadly encompasses the challenge of determining a global position within a pre-mapped area, especially where GNSS is unreliable. This context implies a significant initial uncertainty in pose estimation. Leveraging deep-feature extraction advancements [136], Sarlin et al. [137] introduced a hierarchical localization approach using a unified CNN architecture that integrates both local geometric features and a global descriptor for precise $6$-DoF localization in vast environments. Yin et al. [54] presented a cross-modality visual localization technique tailored for extensive campus areas, employing cross-domain transfer networks. These networks harness condition-invariant features from visual inputs and learn geometric similarities to LiDAR projections, enhancing long-term navigational robustness in dynamically changing environments. Regarding the localization in a larger-scale urban road, Liu et al. [138] tackled the problem by introducing cross-view matching for large-scale geo-localization. Their method, which incorporates both orientation and geometric data, significantly enhances recall rates in spatial localization.
@@
The Teach-and-Repeat (T&R) framework is an efficient navigation solution for diverse mobile robots [139, 140]. Without the requirement of constructing a precise global map, it has achieved great performance in applications such as long-range navigation and planet exploration. During the teach phase, a robot is manually guided along a specific path to generate a topological map, optionally incorporating local metric data. Subsequently, in the repeat phase, the robot autonomously localizes itself within this map to follow the established route, demonstrating an efficient method for traversing pre-determined paths even in changing environments. Therefore, T&R systems necessitate a robust PR module to guarantee the precise localization at the topological level. Chen et al. [141] introduced the sequence matching scheme for enduring T&R operations. Mattamala et al. [140] proposed to dynamically choose the most informative camera during the repeat phase under a multi-camera configuration, mitigating the impact of sudden PR variations. And PlanNav [142] used PR to limit the number of sub-goal candidates for topological navigation.
@@ V-A2 Opportunities
PR is essential for large-scale and long-term navigation, evolving from mere loop closure detection to a broad spectrum of applications including global positioning, failure recovery, and T&R navigation. As PR technology advances, it is timely to revisit navigation system designs, positioning PR as a central element in modern robotic frameworks to enhance navigation in challenging environments. The evolution of mobile robots introduces new challenges for PR, particularly in environments with repetitive features such as multi-floor buildings and underground parking lots. Recent innovations like the hierarchical 3D scene graph proposed by Hughes et al. [66] provide novel PR solutions. Additionally, PR has the potential to enhance the construction of 3D scene graphs by improving spatial segmentation through place similarity comparisons and facilitating efficient exploration in tasks such as object/image-directed exploration (e.g., finding a chair in an office) [143]. These investigations are crucial for advancing dynamic and long-term navigation strategies in complex environments.
@@ V-B1 Brief Survey
The temporal disparities between capturing satellite images and onboard sensor images often span years. Therefore, the conditional changes are mainly attributed to day-night transitions and seasonal variances such as lighting conditions, changes in vegetation, and snow coverage. Current strategies to mitigate these changes [145, 77] include image transformation and feature matching.
Bhavit et al. [146] explored the efficacy of the Normalized Information Distance for aligning Google Earth (GE) satellite images with unmanned aerial vehicle images, demonstrating its superiority over traditional photometric error measurements in day-night scenarios. Following this, an auto-encoder network was developed to map raw images into an embedding space, enhancing robustness against environmental changes by simplifying optimization and storage processes [145]. For addressing seasonal variations, Anthony et al. [144] utilized the U-Net image transform model to align cross-seasonal image pairs. This approach excels at high altitudes, where invariant geometric features prevail across seasons.
However, the majority of previous studies overlooks viewpoint differences (orientation and altitude). iSimLoc [77] utilizes NetVLAD for clustering local features into comprehensive global place descriptors, enhancing feature matching with sequential data. VTRN extends beyond UAVs to ground robots, tackling substantial viewpoint shifts known as “cross-view localization” [138, 147, 148, 149]. An example is visualized in Fig. 4(b). Liu et al. [138] innovated a Siamese network that encodes image pixel orientations, bolstering feature discrimination for localization. SNAP [147] introduced neural representations for Ground Elevation images, dependent solely on ego-view images and camera poses for training, spontaneously generating significant semantics. Shi et al. [148] created a geometry-enhanced cross-view transformer for establishing correspondence across views. In addressing range sensors, Tang et al. [149] developed a representation transforming GE images into 2D point collections for direct comparison with BEV images from LiDAR data.
@@ V-C Multi-Agent Localization and Mapping
Multi-agent systems bring a pivotal shift in addressing complex and dynamic tasks that are beyond the capability of a single agent. The collaboration among robots significantly increase the efficiency to achieve common goals such as cooperative scene exploration [153] and reconstruction [154]. However, one of major challenges in realizing decentralized multi-agent cooperation is to obtain real-time relative coordinates w.r.t. each robot, which become serious in environments characterized by uncertainty and high complexity. PR methods provide a series of solutions, but as pointed out in Section. IV-A and Section. IV-B, the appearance and viewpoint difference from different agents will cause data association to fail for multi-agent cooperation.
@@ V-C1 Brief Survey
Recent advancements in multi-agent systems (MAS) have showcased diverse PR approaches to collaborative mapping and localization. Van et al. [155] introduced a collaborative SLAM system that compresses visual features, allowing for efficient multi-session mapping on the KITTI dataset. Sasaki et al. [156] developed a rover-copter-orbiter cooperative system, leveraging satellite images for coordinated localization among agents and enabling the generation of optimized paths for rover robots through rich textures captured by the copter. Ebadi et al. [157] proposed a geometric-based multi-agent SLAM system for unstable underground environments, employing a robust filtering mechanism to reject noise in data association, demonstrating the critical role of 3D geometric features in the robustness of point-based approaches. Tian et al. [154] introduced the distributed multi-agent metric-semantic SLAM system called Kimera-Multi, where a distributed loop closure detection based on DBoW2 is implemented. Its following work Hydra-Multi [158] employed a multi-robot team to collaboratively construct a 3D scene graph, where scene-graph-based hierarchical loop closure detection is designed. Labbé et al. [159] focused on visual appearance-based loop closure detection methods for multi-session mapping, allowing a single robot to map separate areas across different sessions without requiring initial transformations between trajectories.
These methodologies underscore the evolving landscape of multi-agent localization, setting the groundwork for future cross-disciplinary research [160]. However, challenges remain, particularly in large-scale map merging, where significant perspective and appearance differences pose hurdles. The most recent contribution by Yin et al. [84] addresses these challenges with a framework for large-scale data association and map merging, extracting viewpoint-invariant place descriptors and filtering unreliable loop closures, marking a significant step forward in the field. This body of work not only highlights the potential for multi-agent cooperation but also underscores the necessity of innovative solutions for PR and large-scale data integration in achieving effective map merging and localization.
@@ V-C2 Opportunities
The field of MAS is approaching a period of notable developments, with PR contributing significantly to the evolution of autonomous technologies. Among the most promising avenues is the integration of Neural mapping such as NeRF [10] and 3D Gaussian Splatting [161], offering a groundbreaking approach to rendering photo-realistic and three-dimensional environments from sparse and unstructured data. The PR technique, when applied within systems such as virtual and augmented reality, can enable seamless and immersive interactions between users such as real-world massively multiplayer online games.
Furthermore, PR methods facilitate the applications of MAS in GNSS-denied environments such as subterranean scenes [153], factories, and forests, as demonstrated in the drone swarm system [160]. Exploiting the coordination and communication capabilities of MAS enables safer and more efficient operations in hazardous environments, ranging from deep-sea explorations to space missions and disaster zones, thereby lessening the heavy reliance on communication infrastructure. The collaborative nature of MAS also opens up innovative approaches to crowd-sourcing data collection in robots. Taking the Tesla FSD system [162] as an example, by leveraging a network of vehicles equipped with sensory technologies, a more comprehensive and dynamic mapping of urban roads can be achieved. This provides a large amount of data for algorithm training and thus enhances the safety and reliability of autonomous navigation systems.
@@ V-D1 Brief Survey
Tipaldi et al. [163] introduced a traditional probability-based approach to lifelong localization, leveraging a combination of a particle filter with a hidden Markov model to assess dynamic changes in local maps effectively. Zhao et al. [164] proposed a novel lifelong LiDAR SLAM framework tailored for extended indoor navigation tasks. This framework primarily employs a multiple-session mapping strategy to construct and refine maps while concurrently optimizing memory usage through a Chow-Liu tree-based method [165]. Notably, real-world SLAM implementations tend to struggle more significantly with less dynamic objects, such as parked cars, compared to highly dynamic ones, like moving vehicles. Drawing inspiration from this challenge, Zhu et al. [166] have developed a semantic mapping-enhanced lifelong localization framework that seamlessly integrates existing object detection techniques to update maps continually.
Lifelong place feature learning is crucial for navigation and recognition systems, often challenged by the issue of catastrophic forgetting. This issue arises when systems fail to retain knowledge over long periods, particularly problematic in dynamic or evolving environments. Most existing PR methods focus on short-term or static cases, lacking the ability to adapt and learn continuously without losing previous knowledge. The benchmark on VPR highlights this gap, emphasizing the need for more resilient approaches [167]. An early work to address these challenges was presented by Mactavish et al. [168] A vision-in-the-loop navigation system that incorporates a visual T&R was introduced. This system facilitates long-term, online visual place feature learning and employs a multi-experience localization mechanism. This mechanism assists in matching current observations with relevant past experiences, enhancing the system’s ability to navigate and recognize places over extended periods.
However, real-world applications often involve infinite data streams, leading to scalability challenges. Doan et al. [89] addressed the challenge by integrating a HMM with an innovative two-tiered memory management strategy. This approach effectively segregates active memory from passive storage, facilitating dynamic image transfers essential for lifelong autonomy. Their methodology presents a pragmatic solution for managing long-term navigation tasks, maintaining steady performance without notably elevating time or space demands. Yin et al. [17] proposed BioSLAM, a bio-inspired lifelong learning framework tailored for PR. This innovative framework incorporates a dual-memory system to bolster a robot’s PR capabilities incrementally while mitigating catastrophic forgetting. It encompasses: (1) a dynamic memory for the agile assimilation of new observations; and (2) a static memory to balance fresh insights with established knowledge, guaranteeing consistent PR performance. Furthermore, to evaluate the performance of lifelong systems, two novel evaluation metrics were proposed: adaptation efficiency and retention ability.
@@ VI Datasets & Evaluation
Open datasets introducing new sensor modalities, challenging scenarios, and diverse challenges are instrumental in driving the development of PR approaches. To fairly assess the performance of various PR algorithms and identify their limitations, well-designed evaluation metrics are crucial. In this section, we briefly introduce several public PR datasets, propose a new perspective for evaluation, and discuss open-source libraries relevant to PR. We additionally present our open-source package for general PR research.
@@
  Dataset Scenarios Geographical Coverage Sensors Appearance Diversity Viewpoint Diversity Dynamic Diversity
  Nordland [169] Train ride $748km$ PinC Four seasons
  KITTI [170] Urban Street $1.7km$ L, PinC Day-time
  Oxford RobotCar [171] Urban + Suburban $10km$ L, PinC All kinds
  Mapillary [16] Urban + Suburban $4228km$ PinC All kinds
  KITTI360 [172] Urban Street $73.7km$ L, PinC, PanC Day-time
  ALTO [173](States) Urban+Rural+Nature $50km$ Top-down PinC Day-time
  ALITA [126](City) Urban + Terrain $120km$ L Day time
  ALITA [126](Campus) Campus $60km$ L, PanC Day/Night
  Oxford Radar RoboCar [174] Urban $280km$ L, R, PinC Day/Night, Weather, Traffic
  # L: LiDAR. R: Radar. PinC: Pinhole Camera. PanC: Panoramas Camera.

### v2

@@ IV-D Efficiency
While sequence matching improves localization accuracy, the brute-force approach in SeqSLAM [102] remains computationally intensive, though integrating odometry [149] could enhance performance. Recent efforts to optimize efficiency include particle filters [120], approximate nearest neighbor search [121], dimensionality reduction/quantization [150], Hidden Markov Model (HMM) [122] for sequence modeling, and coarse-to-fine strategies [151, 133, 104] that eliminate exhaustive database searches through candidate sequence initialization.
@@ V-A Long-Term & Large-Scale Navigation
The most in-demand robotic tasks requiring PR is the autonomous navigation, including applications such as autonomous driving [159] and subterranean search [160]. PR enables robots to obtain their global location with precision up to the topological level in a known environment, despite conditional changes of environments. This capability is crucial to safe and reliable navigation, as it allows robots to (1) determine whether this place is visited before, (2) recorver from failure against the kidnapped problem, and (3) progressively update and enhance their navigational maps over time.
@@ V-A1 Brief Survey
SLAM uses PR to correct accumulated drift by recognizing previously visited locations despite environmental changes or viewpoint variations. Systems like ORB-SLAM [161] and VINS-Mono [162] integrate PR modules using the DBoW2 library [25] with BoW-based global descriptors for efficient loop detection, complemented by geometric verification to filter false candidates [163]. LiDAR- and Radar-based SLAM [164, 165] further improve loop closure without requiring an additional camera.
@@
Tracking failure, where the system fails to correctly associate across adjacent frames, often leads to SLAM collapse and can induce the Kidnapped problem. Failures may result from issues such as motion blur, occlusion, or hardware disconnections. Recovery requires relocalizing the robot within the previously built map, posing a more complex PR challenge than simple loop detection, as odometry priors are unavailable. Chen et al. [166] addressed this with a submap-based SLAM system that enhances resilience by creating and integrating submaps upon failure, ensuring mapping continuity with DBoW2. Furthermore, Kuse and Shen [167] optimized [56] for VPR by introducing an all-pair loss function and decoupled convolutions, which accelerates training convergence and reduces the number of parameters. This solution was integrated into a stereo-inertial SLAM system, achieving real-time loop closure detection and reliable failure recovery in complex indoor environments.
@@
Global localization broadly encompasses the challenge of determining a global position within a pre-mapped area, especially where GNSS is unreliable. This context implies a significant initial uncertainty in pose estimation. Sarlin et al. [168] proposed a hierarchical localization method using a unified CNN to combine local geometric features with a global descriptor for precise $6$-DoF localization in large environments. Yin et al. [65] developed a cross-modality visual localization approach for campus-scale areas, leveraging cross-domain transfer networks to align visual inputs with LiDAR projections for robust long-term navigation. For large-scale urban roads, Liu et al. [169] introduced a cross-view matching method integrating orientation and geometric data to improve spatial localization recall rates.
@@
The Teach-and-Repeat (T&R) framework is an efficient navigation solution for diverse mobile robots [170]. Without the requirement of constructing a precise global map, it has achieved great performance in applications such as long-range navigation and planet exploration. During the teach phase, a robot is manually guided along a specific path to generate a topological map, optionally incorporating local metric data. Subsequently, in the repeat phase, the robot autonomously localizes itself within this map to follow the established route, demonstrating an efficient method for traversing pre-determined paths even in changing environments. Therefore, T&R systems necessitate a robust PR module to guarantee the precise localization at the topological level. Chen et al. [171] introduced the sequence matching scheme for enduring T&R operations. Mattamala et al. [170] proposed to dynamically choose the most informative camera during the repeat phase under a multi-camera configuration, mitigating the impact of sudden PR variations. PlaceNav [172] used PR to limit the number of sub-goal candidates for topological navigation.
@@ V-A2 Opportunities
PR is essential for large-scale and long-term navigation, evolving from mere loop closure detection to a broad spectrum of applications including GNSS-denied global localization, failure recovery, and T&R navigation. The field’s maturation drives a paradigm shift of robot navigation, establishing spatial awareness as a primary perceptual layer that directly addresses two key challenges: (1) sustaining large-scale localization accuracy over extended temporal horizons, and (2) enabling dynamic map reuse/updating mechanisms for persistent environmental modeling, significantly enhancing mission success rates and enabling long-horizontal planning.
Recent innovations such as the hierarchical 3D scen graph proposed by Hughes et al. [80] and the text-image retrieval system [90] provide novel PR solutions. A subsequent innovation involves the joint learning of PR and object detection, as proposed in [173], enables scene graph construction by integrating spatial and semantic understanding solely from images. Additionally, PR also facilitates efficient exploration in tasks such as object/image-directed navigation (e.g., finding a chair in an office) [174, 175]. These investigations are crucial for advancing dynamic and long-term navigation strategies in increasingly complex, ever-changing environments.
@@ V-B1 Brief Survey
The temporal disparities between capturing satellite images and onboard sensor images often span years. Therefore, the conditional changes are mainly attributed to day-night transitions and seasonal variances such as lighting conditions, changes in vegetation, and snow coverage. Current strategies to mitigate these changes [177] include image transformation and feature matching.
Bhavit et al. [177] investigated the use of Normalized Information Distance to align Google Earth (GE) satellite images with unmanned aerial vehicle images, showing its advantage over traditional photometric error measures in day-night scenarios. Building on this, an auto-encoder network was introduced to embed raw images, improving robustness to environmental changes and streamlining optimization and storage [178]. To handle seasonal variations, Anthony et al. [176] applied a U-Net image transform model for aligning cross-seasonal images, particularly effective at high altitudes where invariant geometric features dominate across seasons.
Most previous studies have overlooked viewpoint variations, such as differences in orientation and altitude. iSimLoc [100] leverages NetVLAD for local feature aggragation, improving feature matching using sequential data. For ground robots, Sarlin et al. [179] introduced neural representations for Ground Elevation images, relying solely on ego-view images and camera poses, generating rich semantics automatically. Shi et al. [180] developed a geometry-enhanced cross-view transformer for view correspondence, and Tang et al. [181] proposed a method that transforms GE images into 2D point collections to align directly with BEV images from LiDAR data.
@@ V-C Multi-Agent Localization and Mapping
Multi-agent systems bring a pivotal shift in addressing complex and dynamic tasks that are beyond the capability of a single agent. The collaboration among robots significantly increase the efficiency to achieve common goals such as cooperative scene exploration [185]. However, one of major challenges in realizing decentralized multi-agent cooperation is to obtain real-time relative coordinates w.r.t. each robot, which become serious in environments characterized by uncertainty and high complexity. PR methods provide a series of solutions, but as pointed out in Section IV-A and Section IV-B, the appearance and viewpoint difference from different agents will cause data association to fail for multi-agent cooperation.
@@ V-C1 Brief Survey
Recent advancements in multi-agent systems (MAS) have introduced diverse PR strategies for collaborative mapping and localization. Van et al. [186] utilized compresses visual features for efficient multi-session mapping on KITTI. Sasaki et al. [187] developed a rover-copter-orbiter system, using satellite images for coordinated localization and optimized rover paths. Ebadi et al. [188] presented a geometric-based system for unstable environments, employing robust filtering to enhance 3D geometric feature reliability. Kimera-Multi [189] incorporates distributed loop closure detection, while Hydra-Multi [190] enables multi-robot 3D scene graph construction with hierarchical loop closure. Labbé et al. [191] focused on visual LCD, supporting multi-session mapping without initial trajectory transformations. These methodologies underscore the evolving landscape of multi-agent localization, setting the groundwork for future cross-disciplinary research [192]. However, challenges remain, particularly in large-scale map merging, where significant perspective and appearance differences pose hurdles. The most recent contribution by Yin et al. [107] addresses these challenges with a framework for large-scale data association and map merging, extracting viewpoint-invariant place descriptors and filtering unreliable loop closures, marking a significant step forward in the field.
@@ V-C2 Opportunities
The field of MAS is approaching a period of notable developments, with PR contributing significantly to the evolution of autonomous technologies. Among the most promising avenues is the integration of Neural mapping [193, 194] such as 3D Gaussian Splatting [194] that offers a groundbreaking approach to render photo-realistic from a novel view with sparse and unstructured data. The PR technique, when applied within systems such as virtual and augmented reality, can enable seamless and immersive interactions among agents such as real-world massively multi-player online games and human-machine interaction.
Furthermore, PR methods facilitate the applications of MAS in GNSS-denied environments such as subterranean scenes [185], factories, and forests, as demonstrated in the drone swarm system [192]. Leveraging the coordination and communication of MAS enhances safety and efficiency in hazardous environments, from deep-sea and space exploration to disaster response, reducing dependence on communication infrastructure. MAS also enables innovative crowd-sourced data collection, as seen in the Tesla FSD system [195], where a network of sensor-equipped vehicles generates dynamic urban maps, improving algorithm training and autonomous navigation reliability.
@@ V-D1 Brief Survey
Tipaldi et al. [196] introduced a traditional probability-based approach to lifelong localization, leveraging a combination of a particle filter with a HMM to assess dynamic changes in local maps effectively. Zhao et al. [197] proposed a novel lifelong LiDAR SLAM framework tailored for extended indoor navigation tasks. This framework primarily employs a multiple-session mapping strategy to construct and refine maps while concurrently optimizing memory usage through a Chow-Liu tree-based method [198]. Notably, real-world SLAM implementations tend to struggle more significantly with less dynamic objects, such as parked cars, compared to highly dynamic ones, like moving vehicles. Drawing inspiration from this challenge, Zhu et al. [199] have developed a semantic mapping-enhanced lifelong localization framework that seamlessly integrates existing object detection techniques to continuously update maps.
Lifelong feature learning is crucial for navigation systems but faces challenges from catastrophic forgetting, the gradual loss of prior knowledge, particularly in dynamic environments. Most VPR methods operate in short-term or static contexts and struggle with continuous adaptation without performance degradation, as evidenced by VPR benchmarks [200]. Mactavish et al. [201] addressed this with a visual T&R framework, enabling online lifelong feature learning through a multi-experience localization mechanism that dynamically matches current observations to past experiences. Building on this, Chen et al. [171] introduced an experience graph to structurally link temporally disjointed image sequences, facilitating continual data aggregation for network training.
To tackle scalability challenges in lifelong learning systems handling infinite data streams, Doan et al. [112] proposed a hybrid approach combining the HMM with a two-tiered memory architecture. By separating active memory (for real-time processing) from passive storage (for long-term retention), their method enables efficient dynamic image transfer, ensuring stable performance with minimal computational overhead. Yin et al. [16] developed BioSLAM, a lifelong learning framework for VPR. BioSLAM employs a dual-memory system: A dynamic memory to rapidly assimilate new observations, and A static memory to preserve foundational knowledge while integrating novel insights. This design mitigates catastrophic forgetting and maintains consistent VPR accuracy. The authors also introduced two evaluation metrics: adaptation efficiency (speed of learning new data) and retention ability (preservation of prior knowledge), to demonstrate BioSLAM’s superiority over existing methods in incremental learning scenarios.
@@ VI Datasets & Evaluation
Open datasets introducing new sensor modalities, challenging scenarios, and diverse challenges are instrumental in driving the development of PR approaches. To fairly assess the performance of various PR algorithms and identify their limitations, well-designed evaluation metrics are crucial. In this section, we briefly introduce several public PR datasets, propose a new perspective for evaluation, and discuss open-source libraries relevant to PR.
@@
  Dataset Scenarios Length Sensors Appearance Diversity Viewpoint Diversity Dynamic
  Nordland [202] Train ride $748km$ PinC Four seasons No No
  Oxford RobotCar [203] Urban + Suburban $10km$ L, PinC All kinds No Yes
  Mapillary [15] Urban + Suburban $4228km$ PinC All kinds Yes Yes
  KITTI360 [204] Urban Street $73.7km$ L, PinC, PanC Day-time No Yes
  ALTO [205](States) Urban+Rural+Nature $50km$ Top-down PinC Day-time Yes No
  ALITA [157](City) Urban + Terrain $120km$ L Day time Yes Yes
  ALITA [157](Campus) Campus $60km$ L, PanC Day/Night Yes Yes
  Oxford Radar RoboCar [206] Urban $280km$ L, R, PinC Day/Night, Weather, Traffic Yes No
  # L: LiDAR. R: Radar. PinC: Pinhole Camera. PanC: Panoramic Camera.